How many grams of water can be produced if 3.05 L of hydrogen gas @SATP are reacted with
1.48 g of oxygen? Which is the limiting reagent?
1
Expert's answer
2020-05-08T14:05:33-0400

V(H2) = 3.05 L

m(O2) = 1.48 g

2H2 + O2 => 2H2O

n(H2) = V/Vm = 3.05 L/22.4 L/mol = 0.136 mol

n(O2) = 1.48 g/ 32 mol/L = 0.046 mol => limited reagent

n(H2O) = n(O2) = 0.046 mol

m(H2O) = n(H2O)*M(H2O) = 0.046 mol*18 g/mol = 0.828 g